# Generalizing Analogical Inference from Boolean to Continuous Domains

### v1

@@ 1 Introduction
The items $a,b,c,d$ are supposed to be described by a set of related attributes and represented by tuples of attribute values. Attributes may be Boolean, nominal (i.e., finite attribute domains), or real-valued (e.g., embeddings). Modeling analogical reasoning has a long tradition in artificial intelligence, particularly in logic (Aristotle 2011; Prade and Richard 2013), cognitive modeling (Gentner 1983; Hofstadter and Sander 2013), and machine learning (Hu et al. 2023). In recent years, interest in formal analogical inference has grown, driven by its applications in few-shot learning (Hwang, Grauman, and Sha 2013), transfer learning (Badra 2020; Cornuéjols, Murena, and Olivier 2020), and interpretable AI (Hüllermeier 2020). Analogical reasoning has also played a central role in natural language processing, particularly in morphological analysis and word embedding evaluation, where analogies such as “king is to queen as man is to woman” have served as both tasks and benchmarks (Mikolov, Yih, and Zweig 2013; Bouraoui, Jameel, and Schockaert 2018; Gladkova, Drozd, and Matsuoka 2016). More recently, neural models have been designed to detect and explain analogical relations in textual data, including analogies between sentences, concepts, and procedures (Ushio et al. 2021; Jacob, Shani, and Shahaf 2023; Kumar and Schockaert 2023).
@@
Foundational results characterizing the class of functions compatible with Boolean analogical inference were established in (Couceiro et al. 2017), and later works extended analogical classification techniques to nominal and numerical domains (Bounhas and Prade 2023; Couceiro et al. 2020). Theoretical results, notably those of (Couceiro et al. 2018), have established that analogical inference is exact (i.e., error-free) if and only if the labeling function is affine. More generally, they showed that if a Boolean function is $\varepsilon$-close to an affine function, then the probability of making an incorrect analogical prediction is bounded above by $4\varepsilon$, where the probability is taken over the random choice of training sets. This result has been influential in supporting the soundness of analogy-based classifiers in Boolean settings and later extended to a Galois theory of analogical classifiers (Couceiro and Lehtonen 2024).
@@
Despite this progress, two important limitations remain. First, these results are limited to discrete attribute spaces and binary classification tasks. In practice, many real-world applications involve real-valued features and regression tasks, settings that are not covered by current theory. To the best of our knowledge, no prior work has addressed analogical reasoning in the context of regression. Secondly, the theoretical guarantees established in the Boolean setting such as the generalization bound proposed in (Couceiro et al. 2018; Couceiro and Lehtonen 2024) fail to extend to real-valued functions. This limits the application of these results to handle regression.
@@
In this work, we address these shortcomings by revisiting analogical inference from both theoretical and practical perspectives. Firstly, we construct an explicit counterexample demonstrating that the main generalization bound proposed in (Couceiro et al. 2018) fails to hold, even for Boolean functions that are arbitrarily close to affine. This challenges a foundational assumption in the existing theoretical framework. Secondly, we extend analogical inference to real-valued domains by introducing a unified model based on parameterized analogies defined via generalized means. This formulation naturally subsumes classical notions such as additive and geometric analogies and supports analogical reasoning in regression tasks. Thirdly, we introduce functional distances tailored to this generalized setting and derive both uniform and probabilistic bounds on inference error, yielding worst-case and average-case guarantees under smoothness assumptions. Finally, we characterize the class of continuous functions that preserve analogical structures, showing that they correspond to a family of generalized-power functions with well-defined structural properties.
@@ 2.1 Analogical Inference from Boolean to General Cases
In the Boolean case, the most well-known models of analogy are Klein’s model (Klein et al. 1983) of Boolean analogy [^1] Here the columns are precisely the tuples $(a,b,c,d)$ for which the analogy holds. and the so-called minimal model (Miclet and Prade 2009) [^2] Note that $M$ contains only patterns of the form $x:$:y:yx:x::y:y and $x:y::x:y$, for $x,y\in\{0,1\}$. .
@@
In the Boolean case, the problem of finding an $x\in\{0,1\}$ such that $a:b::c:x$ holds, does not always have a solution. For instance, neither $0:1::1:x$ nor $1:0::0:x$ has a solution in the minimal model $M$ (since $0111$, $0110$, $1000$, $1001$ are not valid patterns for an analogical proportion). In fact, a solution exists if and only if $(a\equiv b)\vee(a\equiv c)$ holds. When a solution exists, it is unique and is given by $x=c\equiv(a\equiv b)$ (see also (Lepage 2023)). This corresponds to the original view advocated by S. Klein (Klein 1983), who however applied the latter formula even to the cases $0:1::1:x$ and $1:0::0:x$, where it yields $x=0$ and $x=1$ respectively.
In the nominal case, the situation is similar. The analogical proportion $a:b::c:x$ may have no solution (e.g., in the minimal model, $s:t::t:x$ has no solution as soon as $s\neq t$), and otherwise (if $a=b$ or $a=c$) the solution is unique, and is given by $x=b$ if $a=c$ and $x=c$ if $a=b$. Namely, the solutions of $s:t::s:x$, $s:$:t:xs:s::t:x, and $s:$:s:xs:s::s:x are $x=t$, $x=t$, and $x=s$, respectively. This motivates the following inference pattern, first formalized by (Stroppa and Yvon 2006); see also (Bounhas, Prade, and Richard 2014):
@@
Assuming that the latter four instances are in analogical proportion for each of the $n$ attributes describing them, and that the class labels are known for $\mathbf{a},\mathbf{b},\mathbf{c}$ but unknown for $\mathbf{d}$, then one may infer that the label for $\mathbf{d}$ as a solution of an analogical proportion equation (Bounhas, Prade, and Richard 2017; Couceiro et al. 2017). The effectiveness of this analogical inference rule led to several studies that aimed to determine which classifiers were compatible with the Analogical Inference Principle (AIP) (Miclet, Bayoudh, and Delhay 2008; Bounhas, Prade, and Richard 2017; Couceiro et al. 2017, 2018, 2020; Bounhas and Prade 2024a; Couceiro and Lehtonen 2024).
In the case of Boolean attributes, a key result has been established in (Couceiro et al. 2017), where it was shown that the set of functions for which analogical inference is sound, i.e., no error occurs, are the analogy-preserving (AP) functions, which coincide exactly with the set of affine Boolean functions. More precisely, they showed the following result.
@@
Moreover, when the function is close to being affine, it was also shown that the prediction accuracy remains high (Couceiro et al. 2018). Assuming a uniform distribution and the Hamming distance $d$ on the Boolean function space $\bigcup_{n>0}\mathbb{B}^{\mathbb{B}^{n}}$, the classification performance remained high for classifiers close to the AP functions.
@@
Given a sample set $S\subseteq\mathbb{B}^{n}$ and a function $f\in\mathbb{B}^{\mathbb{B}^{n}}$, the analogical root (Hug et al. 2016) of a given element $\mathbf{x}\in\mathbb{B}^{n}$, denoted by $\mathbf{R}_{S}(\mathbf{x},f)$, is the set
@@ 2.3 Limitations and Drawbacks
When attributes are valued on finite domains $X_{i}$, i.e. nominal case (which includes the Boolean case), the problem of characterizing analogy-preserving functions $f\colon{\bf X}\to X$, for ${\bf X}=X_{1}\times\cdots\times X_{n}$, has been partially addressed for binary classification ($|X|\leq 2$). This has led to the definition of hard AP functions (HAP) $f\colon\mathbf{X}\to X$, which are either “essentially unary” or “quasi-linear”, i.e., for which there exist $\varphi\colon\{0,1\}\to X$ and $\varphi_{i}\colon X_{i}\to\{0,1\}$ ($1\leq i\leq m$) such that $f=\varphi(\varphi_{1}(x_{1})\oplus\dots\oplus\varphi_{n}(x_{n}))$. However, this characterization is limited to nominal attributes, binary labels and constrained models of analogy, typically those satisfying the patterns $(x,x,y,y)$ and $(x,y,x,y)$. In addition, prior work often assumes a decomposable model of analogy, which may not hold in real-world data. These limitations have been critically examined in recent work, such as (Bounhas and Prade 2024a), which revisits analogical proportions and inference under relaxed assumptions.
These limitations motivate the reformulation of classification tasks using analogy-based regression. In the next section, we construct a counterexample that demonstrates the failure of the error bound from (Couceiro et al. 2018), even in the Boolean setting. We will then introduce an analogical inference principle tailored to regression and continuous domains, enabling extensions to multi-class settings. This framework yields new theoretical guarantees under reasonable smoothness and distributional assumptions.
@@ 4 Analogy-Preserving Functions
The generalization of analogical reasoning from Boolean to continuous domains requires a more flexible notion of analogy that can capture numeric structure. To address this, we adopt a parameterized framework based on generalized means, originally studied by Hölder (Hölder 1889), and recently proposed for analogical inference in (Lepage and Couceiro 2024b). This approach defines analogical proportions over real-valued tuples and supports analogical inference in both classification and regression settings.
@@ 4.3 Characterizing Analogy-Preserving Functions
Observe that the underlying domain is the set of nonnegative real numbers. This might sound as a restriction but a good number of applications meet this condition. A large field of application is image processing where the values on the gray channel (see MNIST data for instance (LeCun et al. 1989)) or RGB channels are non-negative real number (sometimes even natural numbers between 0 and 255). Any processing of images involving numerical analogy is thus possible on this kind of data. One may think about image completion, image reconstruction, etc. Another example of field where representations are non-negative real numbers is, at least theoretically, word embedding models. It has been shown that vectors representing words trained from some specific models are concentrated in an orthant of the space, which means that a rotation can make all components in the vectors non-negative (Mimno and Thompson 2017).
@@ 5 Analogy-based Regression
This section explicitly describes the class $AP_{(\mathbf{p};q)}$ of all $(\mathbf{p};q)$-analogy preserving functions under some natural assumptions (see Subsection 5.1), and provides performance guarantees that establish a tight connection between regression errors and distances to the class of analogy preserving functions in various functional spaces (Subsection 5.2). This not only corrects the estimates and performance guarantees provided in (Couceiro et al. 2018), but also generalizes the frameworks of, e.g., (Bounhas, Prade, and Richard 2017; Bounhas and Prade 2023; Bounhas and Prade 2024a; Couceiro et al. 2017, 2018, 2020), to positive reals and to both classification and regression tasks.
@@ 5.2 Performance Guarantees
Inspired by the statement of Theorem 3 that establishes a correspondence between the distance of a Boolean classifier to the class of AP Boolean functions and the probability of that classifier to make classification errors, we seek an analogous (and correct!) result in the setting of analogy based regression. We will make use of background on functional spaces and measure theory and we refer the reader to (Taylor and Kingman 2008) for further background.
@@
To propose an analogue to the probablistic approach proposed in (Couceiro et al. 2017), we will need to assume that $D\in\mathcal{B}$, where $\mathcal{B}$ is the Borel $\sigma$-algebra over $\mathbb{R}_{+}^{n}$, and consider a probability space $(D,\mathcal{B}_{D},\mathbb{P})$, where $\mathcal{B}_{D}$ consists of the elements of $\mathcal{B}$ that contain $D$ (the ideal of $\mathcal{B}$ generated by $D$), and $\mathbb{P}:\mathcal{B}_{D}\to[0,1]$ is a probability measure. For instance, if $D$ is finite, $D\in\mathcal{B}$ and we can naturally choose $\mathbb{P}$ to be the normalized counting measure (i.e., uniform distribution). [^4] However, even in the discrete case, this is not the only setting that could be of interest (for example, we could have a domain with a Poisson distribution), so we will aim to develop a general theory that works for any probability space. This additional structure enables us to define our desired probabilistic distance.
@@ 6 Boolean Case Revisited
We can now attempt to apply the previous results by revisiting the Boolean case. It has been shown that (for the two models of analogy $M$ and $K$), the analogy preserving functions are precisely the affine functions $\mathbb{B}^{n}\to\mathbb{B}$ (see Theorem 1 that was obtained in (Couceiro et al. 2017)).
@@
If the probability measure in use is the normalized counting measure (uniform distribution), $\mathbb{E}(|\iota f-\iota g|)$ is just the (normalized) Hamming distance between $f$ and $g$, used in (Couceiro et al. 2018).
@@ 7 Conclusion
While the focus of this paper is theoretical, aiming to invalidate incorrect claims and establish a sound unifying framework, it also sets the stage for empirical evaluation. Our formulation connects with SOTA analogical models (e.g., (Bounhas and Prade 2024b)) yet extends them to real-valued outputs through a principled analogical regression rule with formal guarantees (Prop. 16–20). This provides the first foundation for implementing analogical regression against methods such as kernel or k-NN regression.
Our framework yields implementable analogical regression rules with closed-form prediction. It also identifies the class of analogy-preserving functions (Th. 12), enabling learnable parametric models. These results provide both a unified theoretical basis and concrete algorithmic formulations for analogical reasoning in regression tasks. These highlighted new avenues for future work. As mentioned earlier, one of the limitations is the underlying assumption that the analogy model is decomposable with respect to each dimension. Despite not being of utmost importance for the current framework, it prevents taking into account synergetic relations between dimensions and thus a full account of emerging concepts in the analogical reasoning process. Also, we have considered a unified model of analogies based on generalized means. However, it would be worthwhile to investigate other recent variants, e.g., based on generalized norms (Prade and Richard 2024).

### v2

@@ 1 Introduction
The items $a,b,c,d$ are supposed to be described by a set of related attributes and represented by tuples of attribute values. Attributes may be Boolean, nominal (i.e., finite attribute domains), or real-valued (e.g., embeddings). Modeling analogical reasoning has a long tradition in artificial intelligence, particularly in logic (aristotle2011nicomachean; prade2013analogical), cognitive modeling (gentner1983structure; hofstadter2013surfaces), and machine learning (hu2023incontext). In recent years, interest in formal analogical inference has grown, driven by its applications in few-shot learning (DBLP:conf/icml/HwangGS13), transfer learning (Badra2020; CornuejolsMO20), and interpretable AI (DBLP:conf/mdai/Hullermeier20). Analogical reasoning has also played a central role in natural language processing, particularly in morphological analysis and word embedding evaluation, where analogies such as “king is to queen as man is to woman” have served as both tasks and benchmarks (mikolov_naacl_13; DBLP:conf/coling/BouraouiJS18; gladkova_naacl_2016). More recently, neural models have been designed to detect and explain analogical relations in textual data, including analogies between sentences, concepts, and procedures (DBLP:conf/acl/UshioASC20; DBLP:conf/emnlp/JacobSS23; DBLP:conf/emnlp/KumarS23).
@@
Foundational results characterizing the class of functions compatible with Boolean analogical inference were established in (CouceiroHPR17), and later works extended analogical classification techniques to nominal and numerical domains (BounhasP23; CouceiroLMPR20). Theoretical results, notably those of (CouceiroHPR18), have established that analogical inference is exact (i.e., error-free) if and only if the labeling function is affine. More generally, they showed that if a Boolean function is $\varepsilon$-close to an affine function, then the probability of making an incorrect analogical prediction is bounded above by $4\varepsilon$, where the probability is taken over the random choice of training sets. This result has been influential in supporting the soundness of analogy-based classifiers in Boolean settings and later extended to a Galois theory of analogical classifiers (Galois2024).
@@
Despite this progress, two important limitations remain. First, these results are limited to discrete attribute spaces and binary classification tasks. In practice, many real-world applications involve real-valued features and regression tasks, settings that are not covered by current theory. To the best of our knowledge, no prior work has addressed analogical reasoning in the context of regression. Secondly, the theoretical guarantees established in the Boolean setting such as the generalization bound proposed in (CouceiroHPR18; Galois2024) fail to extend to real-valued functions. This limits the application of these results to handle regression.
@@
In this work, we address these shortcomings by revisiting analogical inference from both theoretical and practical perspectives. Firstly, we construct an explicit counterexample demonstrating that the main generalization bound proposed in (CouceiroHPR18) fails to hold, even for Boolean functions that are arbitrarily close to affine. This challenges a foundational assumption in the existing theoretical framework. Secondly, we extend analogical inference to real-valued domains by introducing a unified model based on parameterized analogies defined via generalized means. This formulation naturally subsumes classical notions such as additive and geometric analogies and supports analogical reasoning in regression tasks. Thirdly, we introduce functional distances tailored to this generalized setting and derive both uniform and probabilistic bounds on inference error, yielding worst-case and average-case guarantees under smoothness assumptions. Finally, we characterize the class of continuous functions that preserve analogical structures, showing that they correspond to a family of generalized-power functions with well-defined structural properties.
@@ 2.1 Analogical Inference from Boolean to General Cases
In the Boolean case, the most well-known models of analogy are Klein’s model (klein_anthrolopology_1983) of Boolean analogy [^1] Here the columns are precisely the tuples $(a,b,c,d)$ for which the analogy holds. and the so-called minimal model (miclet_sqaru_2009) [^2] Note that $M$ contains only patterns of the form $x:$:y:yx:x::y:y and $x:y::x:y$, for $x,y\in\{0,1\}$. .
@@
In the Boolean case, the problem of finding an $x\in\{0,1\}$ such that $a:b::c:x$ holds, does not always have a solution. For instance, neither $0:1::1:x$ nor $1:0::0:x$ has a solution in the minimal model $M$ (since $0111$, $0110$, $1000$, $1001$ are not valid patterns for an analogical proportion). In fact, a solution exists if and only if $(a\equiv b)\vee(a\equiv c)$ holds. When a solution exists, it is unique and is given by $x=c\equiv(a\equiv b)$ (see also (lepage_iarml_2023)). This corresponds to the original view advocated by S. Klein (Klein1983), who however applied the latter formula even to the cases $0:1::1:x$ and $1:0::0:x$, where it yields $x=0$ and $x=1$ respectively.
In the nominal case, the situation is similar. The analogical proportion $a:b::c:x$ may have no solution (e.g., in the minimal model, $s:t::t:x$ has no solution as soon as $s\neq t$), and otherwise (if $a=b$ or $a=c$) the solution is unique, and is given by $x=b$ if $a=c$ and $x=c$ if $a=b$. Namely, the solutions of $s:t::s:x$, $s:$:t:xs:s::t:x, and $s:$:s:xs:s::s:x are $x=t$, $x=t$, and $x=s$, respectively. This motivates the following inference pattern, first formalized by (SY06); see also (BouPraRicECAI2014):
@@
Assuming that the latter four instances are in analogical proportion for each of the $n$ attributes describing them, and that the class labels are known for $\mathbf{a},\mathbf{b},\mathbf{c}$ but unknown for $\mathbf{d}$, then one may infer that the label for $\mathbf{d}$ as a solution of an analogical proportion equation (BounhasPR17a; CouceiroHPR17). The effectiveness of this analogical inference rule led to several studies that aimed to determine which classifiers were compatible with the Analogical Inference Principle (AIP) (miclet_jair_08; BounhasPR17a; CouceiroHPR17; CouceiroHPR18; CouceiroLMPR20; DBLP:journals/ijar/BounhasP24; Galois2024).
In the case of Boolean attributes, a key result has been established in (CouceiroHPR17), where it was shown that the set of functions for which analogical inference is sound, i.e., no error occurs, are the analogy-preserving (AP) functions, which coincide exactly with the set of affine Boolean functions. More precisely, they showed the following result.
@@
Moreover, when the function is close to being affine, it was also shown that the prediction accuracy remains high (CouceiroHPR18). Assuming a uniform distribution and the Hamming distance $d$ on the Boolean function space $\bigcup_{n>0}\mathbb{B}^{\mathbb{B}^{n}}$, the classification performance remained high for classifiers close to the AP functions.
@@
Given a sample set $S\subseteq\mathbb{B}^{n}$ and a function $f\in\mathbb{B}^{\mathbb{B}^{n}}$, the analogical root (HugPraRicSerECAI2016) of a given element $\mathbf{x}\in\mathbb{B}^{n}$, denoted by $\mathbf{R}_{S}(\mathbf{x},f)$, is the set
@@ 2.3 Limitations and Drawbacks
When attributes are valued on finite domains $X_{i}$, i.e. nominal case (which includes the Boolean case), the problem of characterizing analogy-preserving functions $f\colon{\bf X}\to X$, for ${\bf X}=X_{1}\times\cdots\times X_{n}$, has been partially addressed for binary classification ($|X|\leq 2$). This has led to the definition of hard AP functions (HAP) $f\colon\mathbf{X}\to X$, which are either “essentially unary” or “quasi-linear”, i.e., for which there exist $\varphi\colon\{0,1\}\to X$ and $\varphi_{i}\colon X_{i}\to\{0,1\}$ ($1\leq i\leq m$) such that $f=\varphi(\varphi_{1}(x_{1})\oplus\dots\oplus\varphi_{n}(x_{n}))$. However, this characterization is limited to nominal attributes, binary labels and constrained models of analogy, typically those satisfying the patterns $(x,x,y,y)$ and $(x,y,x,y)$. In addition, prior work often assumes a decomposable model of analogy, which may not hold in real-world data. These limitations have been critically examined in recent work, such as (DBLP:journals/ijar/BounhasP24), which revisits analogical proportions and inference under relaxed assumptions.
These limitations motivate the reformulation of classification tasks using analogy-based regression. In the next section, we construct a counterexample that demonstrates the failure of the error bound from (CouceiroHPR18), even in the Boolean setting. We will then introduce an analogical inference principle tailored to regression and continuous domains, enabling extensions to multi-class settings. This framework yields new theoretical guarantees under reasonable smoothness and distributional assumptions.
@@ 4 Analogy-Preserving Functions
The generalization of analogical reasoning from Boolean to continuous domains requires a more flexible notion of analogy that can capture numeric structure. To address this, we adopt a parameterized framework based on generalized means, originally studied by Hölder (hoelder_mittel_1889), and recently proposed for analogical inference in (lepage_iarml_2024). This approach defines analogical proportions over real-valued tuples and supports analogical inference in both classification and regression settings.
@@ 4.3 Characterizing Analogy-Preserving Functions
Observe that the underlying domain is the set of nonnegative real numbers. This might sound as a restriction but a good number of applications meet this condition. A large field of application is image processing where the values on the gray channel (see MNIST data for instance (lecun_backpropagation_1989)) or RGB channels are non-negative real number (sometimes even natural numbers between 0 and 255). Any processing of images involving numerical analogy is thus possible on this kind of data. One may think about image completion, image reconstruction, etc. Another example of field where representations are non-negative real numbers is, at least theoretically, word embedding models. It has been shown that vectors representing words trained from some specific models are concentrated in an orthant of the space, which means that a rotation can make all components in the vectors non-negative (mimno_emnlp_2017).
@@ 5 Analogy-based Regression
This section explicitly describes the class $AP_{(\mathbf{p};q)}$ of all $(\mathbf{p};q)$-analogy preserving functions under some natural assumptions (see Subsection 5.1), and provides performance guarantees that establish a tight connection between regression errors and distances to the class of analogy preserving functions in various functional spaces (Subsection 5.2). This not only corrects the estimates and performance guarantees provided in (CouceiroHPR18), but also generalizes the frameworks of, e.g., (BounhasPR17a; BounhasP23; DBLP:journals/ijar/BounhasP24; CouceiroHPR17; CouceiroHPR18; CouceiroLMPR20), to positive reals and to both classification and regression tasks.
@@ 5.2 Performance Guarantees
Inspired by the statement of Theorem 3 that establishes a correspondence between the distance of a Boolean classifier to the class of AP Boolean functions and the probability of that classifier to make classification errors, we seek an analogous (and correct!) result in the setting of analogy based regression. We will make use of background on functional spaces and measure theory and we refer the reader to (taylor2008) for further background.
@@
To propose an analogue to the probablistic approach proposed in (CouceiroHPR17), we will need to assume that $D\in\mathcal{B}$, where $\mathcal{B}$ is the Borel $\sigma$-algebra over $\mathbb{R}_{+}^{n}$, and consider a probability space $(D,\mathcal{B}_{D},\mathbb{P})$, where $\mathcal{B}_{D}$ consists of the elements of $\mathcal{B}$ that contain $D$ (the ideal of $\mathcal{B}$ generated by $D$), and $\mathbb{P}:\mathcal{B}_{D}\to[0,1]$ is a probability measure. For instance, if $D$ is finite, $D\in\mathcal{B}$ and we can naturally choose $\mathbb{P}$ to be the normalized counting measure (i.e., uniform distribution). [^4] However, even in the discrete case, this is not the only setting that could be of interest (for example, we could have a domain with a Poisson distribution), so we will aim to develop a general theory that works for any probability space. This additional structure enables us to define our desired probabilistic distance.
@@ 6 Boolean Case Revisited
We can now attempt to apply the previous results by revisiting the Boolean case. It has been shown that (for the two models of analogy $M$ and $K$), the analogy preserving functions are precisely the affine functions $\mathbb{B}^{n}\to\mathbb{B}$ (see Theorem 1 that was obtained in (CouceiroHPR17)).
@@
If the probability measure in use is the normalized counting measure (uniform distribution), $\mathbb{E}(|\iota f-\iota g|)$ is just the (normalized) Hamming distance between $f$ and $g$, used in (CouceiroHPR18).
@@ 7 Conclusion
While the focus of this paper is theoretical, aiming to invalidate incorrect claims and establish a sound unifying framework, it also sets the stage for empirical evaluation. Our formulation connects with SOTA analogical models (e.g., (BounhasP24)) yet extends them to real-valued outputs through a principled analogical regression rule with formal guarantees (Prop. 16–20). This provides the first foundation for implementing analogical regression against methods such as kernel or k-NN regression.
Our framework yields implementable analogical regression rules with closed-form prediction. It also identifies the class of analogy-preserving functions (Th. 12), enabling learnable parametric models. These results provide both a unified theoretical basis and concrete algorithmic formulations for analogical reasoning in regression tasks. These highlighted new avenues for future work. As mentioned earlier, one of the limitations is the underlying assumption that the analogy model is decomposable with respect to each dimension. Despite not being of utmost importance for the current framework, it prevents taking into account synergetic relations between dimensions and thus a full account of emerging concepts in the analogical reasoning process. Also, we have considered a unified model of analogies based on generalized means. However, it would be worthwhile to investigate other recent variants, e.g., based on generalized norms (PradeR24).